# Latent Refinement Decoding: Enhancing Diffusion-Based Language Models by Refining Belief States

### v1

@@ 2 Preliminary
The reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ aims to reconstruct the original data by iteratively denoising from $\mathbf{x}_{T}$ (fully masked) to $\mathbf{x}_{0}$ (clean text). At each denoising step $t$, the model predicts a distribution over tokens for each position: $\hat{p}_{\theta}^{(i)}(\mathbf{x}_{0}|\mathbf{x}_{t})$ for position ii.
@@
where $H_{t}^{(j)}=-\sum_{v}\hat{p}_{\theta}^{(j)}(v|\mathbf{x}_{t})\log\hat{p}_{\theta}^{(j)}(v|\mathbf{x}_{t})$ is the entropy at position $j$, and top-11 selects the position with lowest entropy (highest confidence). This creates a binary embedding assignment: each position uses either $\mathbf{e}_{[\text{MASK}]}$ or a specific token embedding $\mathbf{e}_{v_{t}^{(i)}}$, resulting in complete information loss for positions not selected. This binary decision mechanism creates a discontinuous mapping from probability distributions to discrete embeddings: positions below the confidence threshold are reset to pure [MASK] embeddings, completely discarding their distributional information $\hat{p}_{\theta}(\cdot|\mathbf{x}_{t})$, resulting in abrupt information loss and suboptimal exploration of the posterior distribution.
@@ 3.1 Soft Diffusion
Latent Refinement Decoding mitigates both issues. First, $\hat{q}_{\mathrm{soft}}$ assigns non-zero probability to all tokens, ensuring the true token retains a positive mass even under misprediction. Second, the weighted mixture $\sum_{v}\bar{p}_{t}^{(i)}(v)\mathbf{e}_{v}$ can be viewed as the expected embedding under the model’s belief at position ii. Since self-attention is linear in the embeddings, this representation propagates uncertainty information across positions, enabling different tokens to condition on each other’s belief states.
@@ Appendix B Derivation of the True Posterior in the Masking Process
  • The probability of $x_{t-1}^{(i)}=\texttt{[MASK]}$ is $1-\alpha_{t-1}^{*}$, and once masked, the token remains masked with probability 11. Hence the joint probability is $1-\alpha_{t-1}^{*}$.
@@ Appendix D Stability Analysis of Mixed Embedding Updates
where $\mathbf{e}_{\texttt{[MASK]}}$ denotes the [MASK] embedding, $\mathbf{e}_{v}$ denotes the embedding of token $v$, $\mathcal{T}_{t}^{(i)}$ is the top-$p$ nucleus set at position $i$, and $\bar{p}_{t+1}^{(i)}(v)$ is the renormalised predicted distribution over the nucleus set at position ii.

### v2

@@ 2 Preliminary
The reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ aims to reconstruct the original data by iteratively denoising from $\mathbf{x}_{T}$ (fully masked) to $\mathbf{x}_{0}$ (clean text). At each denoising step $t$, the model predicts a distribution over tokens for each position: $\hat{p}_{\theta}^{(i)}(\mathbf{x}_{0}|\mathbf{x}_{t})$ for position $i$.
@@
where $H_{t}^{(j)}=-\sum_{v}\hat{p}_{\theta}^{(j)}(v|\mathbf{x}_{t})\log\hat{p}_{\theta}^{(j)}(v|\mathbf{x}_{t})$ is the entropy at position $j$, and top-$1$ selects the position with lowest entropy (highest confidence). This creates a binary embedding assignment: each position uses either $\mathbf{e}_{[\text{MASK}]}$ or a specific token embedding $\mathbf{e}_{v_{t}^{(i)}}$, resulting in complete information loss for positions not selected. This binary decision mechanism creates a discontinuous mapping from probability distributions to discrete embeddings: positions below the confidence threshold are reset to pure [MASK] embeddings, completely discarding their distributional information $\hat{p}_{\theta}(\cdot|\mathbf{x}_{t})$, resulting in abrupt information loss and suboptimal exploration of the posterior distribution.
@@ 3.1 Soft Diffusion
Latent Refinement Decoding mitigates both issues. First, $\hat{q}_{\mathrm{soft}}$ assigns non-zero probability to all tokens, ensuring the true token retains a positive mass even under misprediction. Second, the weighted mixture $\sum_{v}\bar{p}_{t}^{(i)}(v)\mathbf{e}_{v}$ can be viewed as the expected embedding under the model’s belief at position $i$. Since self-attention is linear in the embeddings, this representation propagates uncertainty information across positions, enabling different tokens to condition on each other’s belief states.
@@ Appendix B Derivation of the True Posterior in the Masking Process
  • The probability of $x_{t-1}^{(i)}=\texttt{[MASK]}$ is $1-\alpha_{t-1}^{*}$, and once masked, the token remains masked with probability $1$. Hence the joint probability is $1-\alpha_{t-1}^{*}$.
@@ Appendix D Stability Analysis of Mixed Embedding Updates
where $\mathbf{e}_{\texttt{[MASK]}}$ denotes the [MASK] embedding, $\mathbf{e}_{v}$ denotes the embedding of token $v$, $\mathcal{T}_{t}^{(i)}$ is the top-$p$ nucleus set at position $i$, and $\bar{p}_{t+1}^{(i)}(v)$ is the renormalised predicted distribution over the nucleus set at position $i$.